OpenStudy (anonymous):

can somebody please help me on this problem, i am not sure how to work out this problem since it has fractions in it, i have to figure out if it is no solution or no identity

OpenStudy (anonymous):

\[\frac{ 1 }{ 7 }\left( 14-7p \right)-2=-2 (\frac{ 1 }{ 2 }p+3)+6\]

jimthompson5910 (jim_thompson5910):

If you distribute, the fractions will go away \[\Large \frac{ 1 }{ 7 }\left( 14-7p \right)-2=-2 (\frac{ 1 }{ 2 }p+3)+6\] \[\Large \frac{ 1 }{ 7 }*14+\frac{ 1 }{ 7 }(-7p)-2=-2*\frac{ 1 }{ 2 }p-2*3+6\] \[\Large \frac{ 14 }{ 7 }+\frac{ -7p }{ 7 }-2=\frac{ -2 }{ 2 }p-6+6\] \[\Large 2-p-2=-p-6+6\] I'll let you finish

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw